# The life of central radio galaxies in clusters: AGN-ICM studies of eRASS1 clusters in the ASKAP fields

### v1

@@ Appendix A X-ray and radio properties of the eRASS1/ASKAP clusters
[…]
J211652.8-593039 319.2222 -59.5088 0.0578 $60.74\pm 1.36$ $226.0\pm 9.0$ PS1
[…]
J213505.3-625454 323.7732 -62.9159 0.2228 $4.26\pm 0.23$ $298.0\pm 10.2$ PS1
J213536.8-572622 323.9060 -57.4419 0.4268 $4.63\pm 0.24$ $258.0\pm 9.5$ PS1
J213800.9-600758 324.5035 -60.1317 0.3188 $5.88\pm 0.26$ $504.0\pm 13.8$ PS1
[…]
J214553.0-564448 326.4664 -56.7483 0.4810 $2.76\pm 0.20$ $432.0\pm 12.6$ PS1
[…]
J214844.5-611650 327.1785 -61.2795 0.5722 $2.54\pm 0.19$ $364.0\pm 11.4$ PS1
J215129.7-552019 327.8746 -55.3369 0.0409 $2623.45\pm 52.61$ $82.0\pm 6.5$ PS1
[…]
J220153.8-595644 330.4722 -59.9454 0.1066 $2.13\pm 0.19$ $160.0\pm 7.8$ PS1
J220920.7-514811 332.3402 -51.8072 0.1169 $0.79\pm 0.16$ $232.0\pm 9.1$ PS1
[…]
J221504.0-520501 333.7622 -52.0809 0.5002 $2.12\pm 0.18$ $96.0\pm 6.7$ PS1
J221631.8-532501 334.1310 -53.4125 0.1801 $83.86\pm 1.82$ $80.0\pm 6.4$ PS1
J222251.8-483423 335.7113 -48.5765 0.6517 $0.48\pm 0.15$ $166.0\pm 7.9$ PS1
J203328.8-593552 308.3800 -59.5971 0.2004 $353.54\pm 7.21$ $296.0\pm 10.2$ PS1
[…]
J202321.6-553524 305.8370 -55.5972 0.2305 $8.55\pm 0.74$ $238.0\pm 9.2$ PS1
[…]
J203220.5-562738 308.0586 -56.4368 0.2844 $48.45\pm 0.82$ $82.5\pm 4.9$ PS1
J203601.2-513931 308.9892 -51.6527 0.2719 $111.78\pm 1.64$ $286.0\pm 7.6$ PS1
[…]
J204408.5-603931 311.0434 -60.6559 0.1213 $1.79\pm 0.61$ $267.2\pm 8.9$ PS1
J204822.8-613113 312.0900 -61.5172 0.1076 $0.30\pm 0.14$ $37.0\pm 18.5$ PS1
[…]
J205943.1-501908 314.9220 -50.3029 0.3307 $0.59\pm 0.11$ $112.3\pm 4.9$ PS1
J210114.4-554134 315.3039 -55.6940 0.2599 $0.10$ $34.7$ PS1
[…]
J210604.1-584425 316.5194 -58.7412 1.1263 $7.73\pm 0.45$ $87.0\pm 4.7$ PS1
[…]
J212433.5-612500 321.1449 -61.4124 0.4365 $17.47\pm 0.57$ $74.9\pm 37.5$ PS1
J212623.2-553221 321.6349 -55.5222 0.8570 $0.70\pm 0.02$ $160.4\pm 12.8$ PS1
[…]
J215406.4-575136 328.5175 -57.8676 0.0753 $1.67\pm 0.59$ $280.6\pm 5.7$ PS1
[…]
J215826.3-602359 329.6091 -60.4261 0.0849 $2.88\pm 0.21$ $281.4\pm 11.7$ PS1
[…]
J220504.3-592716 331.2512 -59.4547 0.3504 $2.14\pm 0.21$ $144.5\pm 14.5$ PS1
[…]
J220731.3-492529 331.8876 -49.4122 0.2541 $13.18\pm 0.64$ $135.2\pm 7.6$ PS1
J220756.6-522333 331.9916 -52.3887 0.1106 $0.24\pm 0.12$ $39.1\pm 2.3$ PS1
[…]
J221117.3-483401 332.8222 -48.5727 0.2565 $0.29\pm 0.20$ $117.3\pm 3.9$ PS1
J221454.0-532017 333.7287 -53.3502 0.3381 $145.74\pm 1.46$ $124.9\pm 13.5$ PS1
J221959.1-581546 334.9924 -58.2725 0.2814 $7.00\pm 0.79$ $119.1\pm 3.3$ PS1
J221959.2-482901 335.0168 -48.4767 0.8711 $2.07\pm 0.31$ $197.8\pm 8.8$ PS1
J222000.0-522730 335.0352 -52.4640 0.1065 $0.55\pm 0.12$ $204.8\pm 10.3$ PS1
J222000.3-620520 335.0116 -62.0862 0.0634 $0.77\pm 0.29$ $84.7\pm 3.2$ PS1
J222120.9-501707 335.3608 -50.2793 0.1791 $605.73\pm 3.79$ $105.3\pm 4.8$ PS1
J222327.0-522740 335.8409 -52.4637 0.2734 $0.46\pm 0.21$ $221.7\pm 8.2$ PS1
[…]
J222319.6-501552 335.8332 -50.2658 0.2659 $145.86\pm 4.40$ $508.9\pm 10.0$ PS1
[…]
J202555.3-511709 306.4823 -51.2745 0.2286 $25.59\pm 3.80$ $214.1\pm 8.9$ PS1
[…]
J215919.0-521025 329.8301 -52.1648 0.4982 $3.77\pm 0.59$ $96.5\pm 4.1$ PS1
J082604.4+041910 126.5118 4.3250 0.4742 $4.52\pm 0.29$ $46.6\pm 2.7$ SWAG-X
[…]
J083205.0+041853 128.0177 4.3099 0.2012 $152.11\pm 0.68$ $349.1\pm 6.2$ SWAG-X
J083652.6+030000 129.2144 3.0004 0.1910 $5.12\pm 0.09$ $203.1\pm 14.5$ SWAG-X
[…]
J085751.0+031014 134.4750 3.1766 0.2025 $8.24\pm 0.25$ $438.3\pm 7.6$ SWAG-X
[…]
J091235.0-042929 138.1475 -4.4918 0.2060 $16.72\pm 0.17$ $275.3\pm 12.8$ SWAG-X
J091414.7+001922 138.5613 0.3252 0.1665 $10.53\pm 0.18$ $302.3\pm 5.1$ SWAG-X
J091453.6+041611 138.7190 4.2654 0.1477 $84.60\pm 0.24$ $349.2\pm 1.2$ SWAG-X
J092023.1+013433 140.0980 1.5789 0.6939 $5.45\pm 0.16$ $115.4\pm 0.5$ SWAG-X
[…]
J093347.9+071429 143.4528 7.2459 0.2527 $0.81\pm 0.09$ $201.4\pm 21.5$ SWAG-X
[…]
J093640.3+064147 144.1691 6.7115 0.2816 $73.06\pm 0.37$ $278.5\pm 13.1$ SWAG-X
[…]
J094023.3+022824 145.1024 2.4776 0.1509 $10.12\pm 0.37$ $123.3\pm 4.2$ SWAG-X
[…]
J083934.2-014035 129.8798 -1.6729 0.2698 $0.16$ $267.9$ SWAG-X
J083948.5+072232 129.9572 7.4120 0.2006 $0.16$ $115.4$ SWAG-X
[…]
J084147.8-031154 130.4492 -3.1946 0.2218 $0.16$ $105.4$ SWAG-X
[…]
J090128.8-013837 135.3778 -1.6548 0.3152 $0.16$ $115.3$ SWAG-X
[…]
J091756.4+050849 139.4822 5.1524 0.5890 $0.39\pm 0.13$ $143.2\pm 0.8$ SWAG-X
J092050.7+024512 140.2071 2.7539 0.2819 $1.86\pm 0.15$ $300.1\pm 0.3$ SWAG-X
[…]
J092210.0+034626 140.5319 3.7664 0.2679 $0.16$ $567.4$ SWAG-X
J092211.2-002812 140.5456 -0.4603 0.3198 $4.85\pm 0.35$ $244.1\pm 21.3$ SWAG-X
[…]
J093521.3+023222 143.8376 2.5432 0.4993 $21.94\pm 0.89$ $143.1\pm 2.0$ SWAG-X
[…]
J093512.7+004735 143.8012 0.8256 0.3570 $15.21\pm 2.31$ $153.1\pm 0.9$ SWAG-X
J093828.0-041933 144.6500 -4.3351 0.0807 $900.81\pm 24.12$ $345.2\pm 0.3$ SWAG-X
J093439.4+054141 143.6624 5.6957 0.5590 $0.24\pm 0.07$ $68.8\pm 4.9$ SWAG-X
[…]
J091608.3-002355 139.0385 -0.4045 0.3234 $25.05\pm 0.79$ $121.0\pm 3.3$ SWAG-X
J085029.4+001453 132.6159 0.2502 0.1955 $81.03\pm 3.04$ $159.8\pm 33.6$ SWAG-X
[…]
J091406.0-034208 138.5319 -3.7364 0.1622 $37.98\pm 0.98$ $59.2\pm 29.6$ SWAG-X
J022201.8-113233 35.5030 -11.5399 0.3256 $1.48\pm 0.54$ $64.9\pm 32.5$ PS2
[…]
J023043.3-105615 37.6769 -10.9365 0.6498 $3.03\pm 0.51$ $56.5\pm 28.3$ PS2
[…]
J023419.1-095138 38.5830 -9.8656 0.6047 $1.25\pm 0.28$ $91.6\pm 20.4$ PS2
[…]
J024106.3-111241 40.2841 -11.2125 0.0961 $5.72\pm 0.38$ $237.8\pm 69.7$ PS2
[…]
J024424.7-073636 41.1028 -7.6059 0.3474 $4.96\pm 0.92$ $223.5\pm 17.1$ PS2

### v2

@@ Appendix A X-ray and radio properties of the eRASS1/ASKAP clusters
J205316.0-620912CC 313.3130 -62.1517 0.3953 $0.31\pm 0.15$ $218.0\pm 8.8$ PS1
J205555.4-545538CC 313.9830 -54.9302 0.1395 $21.86\pm 0.58$ $374.0\pm 11.6$ PS1
J211250.8-531753CC 318.2121 -53.2983 0.2231 $8.62\pm 0.31$ $130.0\pm 7.3$ PS1
J211652.8-593039 319.2222 -59.5088 0.0578 $60.74\pm 1.36$ $64.0\pm 6.1$ PS1
J212023.5-542845CC 320.1055 -54.4780 0.2410 $5.83\pm 0.26$ $298.0\pm 10.2$ PS1
J213305.7-594531CC 323.2733 -59.7603 0.5051 $0.26\pm 0.15$ $504.0\pm 13.8$ PS1
J213505.3-625454 323.7732 -62.9159 0.2228 $4.26\pm 0.23$ $206.0\pm 8.6$ PS1
J213536.8-572622 323.9060 -57.4419 0.4268 $4.63\pm 0.24$ $432.0\pm 12.6$ PS1
J213800.9-600758 324.5035 -60.1317 0.3188 $5.88\pm 0.26$ $170.0\pm 8.0$ PS1
J214359.4-563717CC 325.9968 -56.6225 0.0820 $168.40\pm 3.51$ $96.0\pm 6.7$ PS1
J214553.0-564448 326.4664 -56.7483 0.4810 $2.76\pm 0.20$ $364.0\pm 11.4$ PS1
J214622.3-571714CC 326.5942 -57.2868 0.0733 $4.69\pm 0.24$ $82.0\pm 6.5$ PS1
J214647.5-573648CC 326.7027 -57.6150 0.6109 $1.02\pm 0.16$ $328.0\pm 10.8$ PS1
J214844.5-611650 327.1785 -61.2795 0.5722 $2.54\pm 0.19$ $316.0\pm 10.6$ PS1
J215129.7-552019 327.8746 -55.3369 0.0409 $2623.45\pm 52.61$ $160.0\pm 7.8$ PS1
J215625.0-513110CC 329.1034 -51.5220 0.4942 $1.94\pm 0.18$ $232.0\pm 9.1$ PS1
J220112.5-614737CC 330.3022 -61.7947 0.2382 $10.27\pm 0.35$ $170.0\pm 8.0$ PS1
J220153.8-595644 330.4722 -59.9454 0.1066 $2.13\pm 0.19$ $96.0\pm 6.7$ PS1
J220920.7-514811 332.3402 -51.8072 0.1169 $0.79\pm 0.16$ $80.0\pm 6.4$ PS1
J221033.0-570945CC 332.6328 -57.1649 0.3003 $0.93\pm 0.16$ $166.0\pm 7.9$ PS1
J221504.0-520501 333.7622 -52.0809 0.5002 $2.12\pm 0.18$ $296.0\pm 10.2$ PS1
J221631.8-532501 334.1310 -53.4125 0.1801 $83.86\pm 1.82$ $160.0\pm 7.8$ PS1
J222251.8-483423 335.7113 -48.5765 0.6517 $0.48\pm 0.15$ $238.0\pm 9.2$ PS1
J203328.8-593552 308.3800 -59.5971 0.2004 $353.54\pm 7.21$ $252.0\pm 9.4$ PS1
J201902.9-564238CC 304.7588 -56.7111 0.2313 $23.98\pm 0.60$ $286.0\pm 7.6$ PS1
J202321.6-553524 305.8370 -55.5972 0.2305 $8.55\pm 0.74$ $175.2\pm 7.6$ PS1
J203153.0-535148CC 307.9693 -53.8626 0.1711 $1.11\pm 0.44$ $108.5\pm 6.4$ PS1
J203220.5-562738 308.0586 -56.4368 0.2844 $48.45\pm 0.82$ $267.2\pm 8.9$ PS1
J203601.2-513931 308.9892 -51.6527 0.2719 $111.78\pm 1.64$ $366.3\pm 8.6$ PS1
J203732.2-540710CC 309.3828 -54.1204 0.1314 $4.57\pm 0.44$ $112.3\pm 4.9$ PS1
J203919.4-624401CC 309.7576 -62.7652 0.1014 $0.10$ $34.7$ PS1
J204125.2-551440CC 310.3542 -55.2459 0.2186 $0.72\pm 0.08$ $131.7\pm 7.4$ PS1
J204408.5-603931 311.0434 -60.6559 0.1213 $1.79\pm 0.61$ $87.0\pm 4.7$ PS1
J204822.8-613113 312.0900 -61.5172 0.1076 $0.30\pm 0.14$ $59.0\pm 4.4$ PS1
J205750.2-524548CC 314.4707 -52.7733 0.2559 $6.52\pm 0.31$ $186.7\pm 8.2$ PS1
J205943.1-501908 314.9220 -50.3029 0.3307 $0.59\pm 0.11$ $184.8\pm 10.0$ PS1
J210114.4-554134 315.3039 -55.6940 0.2599 $0.10$ $74.9$ PS1
J210452.0-514939CC 316.2138 -51.8235 0.0506 $1.64\pm 0.24$ $55.9\pm 2.1$ PS1
J210604.1-584425 316.5194 -58.7412 1.1263 $7.73\pm 0.45$ $337.9\pm 17.1$ PS1
J210655.8-603441CC 316.7334 -60.5756 0.1596 $40.98\pm 0.73$ $280.6\pm 5.7$ PS1
J211032.5-584227CC 317.6634 -58.7150 0.3555 $0.10$ $93.0$ PS1
J212251.3-582948CC 320.7058 -58.4950 0.2932 $0.93\pm 0.16$ $206.0\pm 9.2$ PS1
J212433.5-612500 321.1449 -61.4124 0.4365 $17.47\pm 0.57$ $281.4\pm 11.7$ PS1
J212623.2-553221 321.6349 -55.5222 0.8570 $0.70\pm 0.02$ $287.0\pm 15.9$ PS1
J213151.8-500345CC 322.9636 -50.0624 0.4570 $0.12\pm 0.09$ $144.5\pm 14.5$ PS1
J214528.9-513623CC 326.3730 -51.6071 0.0538 $0.42\pm 0.16$ $39.1\pm 2.3$ PS1
J215406.4-575136 328.5175 -57.8676 0.0753 $1.67\pm 0.59$ $84.2\pm 3.1$ PS1
J215453.8-611541CC 328.7207 -61.2607 0.1033 $1.44\pm 0.92 11$$17.3\pm 3.9$ PS1
J215815.7-502328CC 329.5744 -50.3977 0.4885 $0.11\pm 0.05$ $124.9\pm 13.5$ PS1
J215826.3-602359 329.6091 -60.4261 0.0849 $2.88\pm 0.21$ $119.1\pm 3.3$ PS1
J215918.0-564510CC 329.8259 -56.7500 0.2794 $1.03\pm 0.09$ $197.8\pm 8.8$ PS1
J220504.3-592716 331.2512 -59.4547 0.3504 $2.14\pm 0.21$ $204.8\pm 10.3$ PS1
J220609.1-494126CC 331.5336 -49.7188 0.1284 $5.42\pm 0.42$ $105.3\pm 4.8$ PS1
J220731.3-492529 331.8876 -49.4122 0.2541 $13.18\pm 0.64$ $221.7\pm 8.2$ PS1
J220756.6-522333 331.9916 -52.3887 0.1106 $0.24\pm 0.12$ $50.0\pm 4.6$ PS1
J220952.8-553521CC 332.4480 -55.5817 0.1680 $135.89\pm 2.37$ $312.8\pm 5.9$ PS1
J221117.3-483401 332.8222 -48.5727 0.2565 $0.29\pm 0.20$ $148.4\pm 9.9$ PS1
J221454.0-532017 333.7287 -53.3502 0.3381 $145.74\pm 1.46$ $508.9\pm 10.0$ PS1
J221959.1-581546 334.9924 -58.2725 0.2814 $7.00\pm 0.79$ $214.1\pm 8.9$ PS1
J221959.2-482901 335.0168 -48.4767 0.8711 $2.07\pm 0.31$ $378.2\pm 16.2$ PS1
J222000.0-522730 335.0352 -52.4640 0.1065 $0.55\pm 0.12$ $96.5\pm 4.1$ PS1
J222000.3-620520 335.0116 -62.0862 0.0634 $0.77\pm 0.29$ $46.6\pm 2.7$ PS1
J222120.9-501707 335.3608 -50.2793 0.1791 $605.73\pm 3.79$ $349.1\pm 6.2$ PS1
J222327.0-522740 335.8409 -52.4637 0.2734 $0.46\pm 0.21$ $129.6\pm 9.4$ PS1
J222420.4-503901CC 336.0797 -50.6577 0.3354 $0.10$ $89.5$ PS1
J211622.6-595403CC 319.0925 -59.9032 0.0602 $101.26\pm 6.48$ $351.2\pm 2.4$ PS1
J213917.7-630845CC 324.8008 -63.1455 0.2373 $0.34\pm 0.20$ $105.1\pm 8.9$ PS1
J222319.6-501552 335.8332 -50.2658 0.2659 $145.86\pm 4.40$ $552.6\pm 8.5$ PS1
J201847.9-524238CC 304.6946 -52.6910 0.0502 $309.06\pm 5.06$ $181.6\pm 2.0$ PS1
J202555.3-511709 306.4823 -51.2745 0.2286 $25.59\pm 3.80$ $438.3\pm 7.6$ PS1
J220052.4-515502CC 330.2247 -51.9221 0.2180 $6.32\pm 0.36$ $209.5\pm 7.3$ PS1
J215919.0-521025 329.8301 -52.1648 0.4982 $3.77\pm 0.59$ $275.3\pm 12.8$ PS1
J082604.4+041910 126.5118 4.3250 0.4742 $4.52\pm 0.29$ $302.3\pm 5.1$ SWAG-X
J083133.9-035152CC 127.8926 -3.8652 0.1746 $2.64\pm 0.16 11$$15.4\pm 0.5$ SWAG-X
J083205.0+041853 128.0177 4.3099 0.2012 $152.11\pm 0.68$ $271.3\pm 0.2$ SWAG-X
J083652.6+030000 129.2144 3.0004 0.1910 $5.12\pm 0.09$ $116.8\pm 0.7$ SWAG-X
J083714.1-032144CC 129.3078 -3.3617 0.0410 $0.40\pm 0.09$ $30.3\pm 2.1$ SWAG-X
J083811.9-015938CC 129.5491 -1.9932 0.5590 $75.68\pm 0.31$ $242.3\pm 0.1$ SWAG-X
J084306.7+002834CC 130.7832 0.4831 0.2636 $10.24\pm 0.10$ $154.6\pm 0.5$ SWAG-X
J084342.8+040323CC 130.9295 4.0596 0.2021 $0.69\pm 0.07$ $123.3\pm 4.2$ SWAG-X
J084531.1+022840CC 131.3825 2.4759 0.0768 $10.93\pm 0.14$ $68.3\pm 0.2$ SWAG-X
J085412.1-022058CC 133.5539 -2.3495 0.3684 $0.63\pm 0.13$ $258.0\pm 21.1$ SWAG-X
J085751.0+031014 134.4750 3.1766 0.2025 $8.24\pm 0.25$ $267.9\pm 3.3$ SWAG-X
J090131.4+030055CC 135.3795 3.0157 0.1936 $0.36\pm 0.09$ $115.4\pm 9.1$ SWAG-X
J091235.0-042929 138.1475 -4.4918 0.2060 $16.72\pm 0.17$ $124.8\pm 0.2$ SWAG-X
J091414.7+001922 138.5613 0.3252 0.1665 $10.53\pm 0.18$ $138.1\pm 0.9$ SWAG-X
J091453.6+041611 138.7190 4.2654 0.1477 $84.60\pm 0.24$ $115.3\pm 0.1$ SWAG-X
J092023.1+013433 140.0980 1.5789 0.6939 $5.45\pm 0.16$ $337.9\pm 3.8$ SWAG-X
J092546.2-014321CC 141.4387 -1.7288 0.2265 $95.17\pm 0.56$ $300.1\pm 0.3$ SWAG-X
J092647.6+050038CC 141.6970 5.0011 0.4659 $2.09\pm 0.18$ $567.4\pm 50.9$ SWAG-X
J093150.9-002220CC 142.9627 -0.3705 0.3353 $0.59\pm 0.15$ $280.8\pm 27.9$ SWAG-X
J093346.9-025503CC 143.4550 -2.9139 0.1290 $4.07\pm 0.22$ $143.7\pm 2.9$ SWAG-X
J093347.9+071429 143.4528 7.2459 0.2527 $0.81\pm 0.09$ $137.1\pm 4.6$ SWAG-X
J093527.1+073020CC 143.8644 7.5073 0.2150 $8.18\pm 0.22$ $153.1\pm 0.9$ SWAG-X
J093640.3+064147 144.1691 6.7115 0.2816 $73.06\pm 0.37$ $345.2\pm 0.3$ SWAG-X
J093831.3-012524CC 144.6255 -1.4258 0.4094 $27.17\pm 0.18$ $252.1\pm 0.4$ SWAG-X
J094023.3+022824 145.1024 2.4776 0.1509 $10.12\pm 0.37$ $211.5\pm 2.4$ SWAG-X
J083102.8-041145CC 127.7634 -4.1959 0.0334 $79.36\pm 3.56$ $121.0\pm 3.3$ SWAG-X
J083257.5-022735CC 128.2403 -2.4601 0.7484 $0.23\pm 0.04$ $159.8\pm 33.6$ SWAG-X
J083934.2-014035 129.8798 -1.6729 0.2698 $0.16$ $74.0$ SWAG-X
J083948.5+072232 129.9572 7.4120 0.2006 $0.16$ $59.2$ SWAG-X
J084015.0-032654CC 130.0625 -3.4478 0.2261 $0.16$ $64.9$ SWAG-X
J084147.8-031154 130.4492 -3.1946 0.2218 $0.16$ $64.0$ SWAG-X
J084527.7+032736CC 131.3657 3.4608 0.3336 $0.16$ $85.7$ SWAG-X
J084641.8-034334CC 131.6886 -3.7410 0.1892 $0.16$ $56.5$ SWAG-X
J085217.5-010126CC 133.0695 -1.0266 0.4589 $0.16$ $104.4$ SWAG-X
J085642.7-025957CC 134.1780 -2.9934 0.2313 $0.31\pm 0.07$ $91.6\pm 20.4$ SWAG-X
J090128.8-013837 135.3778 -1.6548 0.3152 $0.16$ $82.5$ SWAG-X
J090539.7+043433CC 136.4161 4.5775 0.2277 $0.16$ $65.3$ SWAG-X
J091214.5-021729CC 138.0637 -2.2944 0.1592 $0.16$ $49.1$ SWAG-X
J091756.4+050849 139.4822 5.1524 0.5890 $0.39\pm 0.13$ $237.8\pm 69.7$ SWAG-X
J092050.7+024512 140.2071 2.7539 0.2819 $1.86\pm 0.15$ $152.7\pm 16.4$ SWAG-X
J092121.0+031735CC 140.3380 3.2870 0.3439 $5.70\pm 0.27$ $223.5\pm 17.1$ SWAG-X
J092210.0+034626 140.5319 3.7664 0.2679 $0.16$ $73.6$ SWAG-X
J092211.2-002812 140.5456 -0.4603 0.3198 $4.85\pm 0.35$ $192.7\pm 18.8$ SWAG-X
J092339.7+052647CC 140.9125 5.4467 0.3760 $20.55\pm 0.48$ $246.4\pm 16.2$ SWAG-X
J093025.3+021707CC 142.6027 2.2903 0.5399 $1.20\pm 0.17$ $227.9\pm 33.9$ SWAG-X
J093358.7+071512CC 143.4932 7.2531 0.2526 $37.96\pm 1.08$ $276.6\pm 13.1$ SWAG-X
J093459.9+005438CC 143.7528 0.9041 0.3610 $0.16$ $90.3$ SWAG-X
J093521.3+023222 143.8376 2.5432 0.4993 $21.94\pm 0.89$ $471.6\pm 23.8$ SWAG-X
J085435.8+003858CC 133.6525 0.6426 0.1073 $61.60\pm 1.91$ $152.8\pm 7.2$ SWAG-X
J093512.7+004735 143.8012 0.8256 0.3570 $15.21\pm 2.31$ $692.2\pm 71.4$ SWAG-X
J093828.0-041933 144.6500 -4.3351 0.0807 $900.81\pm 24.12$ $576.1\pm 7.0$ SWAG-X
J093439.4+054141 143.6624 5.6957 0.5590 $0.24\pm 0.07$ $160.8\pm 40.5$ SWAG-X
J091315.0+034847CC 138.3116 3.8136 0.4561 $14.39\pm 0.42$ $282.6\pm 18.7$ SWAG-X
J091510.8+051440CC 138.7951 5.2418 0.1361 $13.72\pm 0.45$ $153.4\pm 8.0$ SWAG-X
J085932.4+030832CC 134.8847 3.1447 0.1955 $0.58\pm 0.12$ $116.0\pm 20.3$ SWAG-X
J091608.3-002355 139.0385 -0.4045 0.3234 $25.05\pm 0.79$ $269.0\pm 15.5$ SWAG-X
J085029.4+001453 132.6159 0.2502 0.1955 $81.03\pm 3.04$ $487.3\pm 14.6$ SWAG-X
J082809.6-000958CC 127.0383 -0.1665 0.0805 $0.34\pm 0.07$ $36.1\pm 7.7$ SWAG-X
J091406.0-034208 138.5319 -3.7364 0.1622 $37.98\pm 0.98$ $218.8\pm 9.1$ SWAG-X
J022201.8-113233 35.5030 -11.5399 0.3256 $1.48\pm 0.54$ $353.5\pm 10.3$ PS2
J022813.4-115703CC 37.0569 -11.9541 0.2003 $1.86\pm 0.62$ $250.4\pm 7.2$ PS2
J022940.1-121406CC 37.4136 -12.2356 0.2760 $0.98\pm 0.54$ $521.7\pm 9.8$ PS2
J023043.3-105615 37.6769 -10.9365 0.6498 $3.03\pm 0.51$ $390.4\pm 14.6$ PS2
J023337.6-094238CC 38.4123 -9.7154 0.2629 $20.45\pm 1.22$ $269.9\pm 8.4$ PS2
J023419.1-095138 38.5830 -9.8656 0.6047 $1.25\pm 0.28$ $418.9\pm 14.2$ PS2
J023419.6-094007CC 38.5798 -9.6676 0.8699 $2.13\pm 0.54$ $437.4\pm 16.5$ PS2
J023720.6-080141CC 39.3375 -8.0473 0.2106 $3.82\pm 0.51$ $232.1\pm 7.2$ PS2
J024037.9-120201CC 40.1725 -12.0638 0.5204 $0.22$ $116.3$ PS2
J024106.3-111241 40.2841 -11.2125 0.0961 $5.72\pm 0.38$ $88.7\pm 3.7$ PS2
J024305.2-093501CC 40.7658 -9.5823 0.1558 $295.23\pm 0.55$ $398.2\pm 5.6$ PS2
J024424.7-073636 41.1028 -7.6059 0.3474 $4.96\pm 0.92$ $261.4\pm 10.3$ PS2